OpenStudy (anonymous):

A plant fertilizer contains 15% by mass nitrogen (N). In a container of soluble plant food, there are 10.5 oz of fertilizer. How many grams of nitrogen are in the container?

OpenStudy (danjs):

15% of the total mass is Nitrogen. Mass of the Nitrogen is - \[0.15*10.5 ~oz = 1.575~oz~Nitrogen\]

OpenStudy (danjs):

1 oz. = 28.35 grams \[\frac{ 1.575 ~oz ~N}{ 1 }*\frac{ 28.35~g~N }{ 1~oz~N } \approx 44.65~g~Nitrogen\]